The sum of two numbers is 84. The second is 9 more than 4 times the first. What are the two numbers?
Answer by Nate(3500)

first number = x
second number = 4x + 9
(x) + (4x + 9) = 84
5x + 9 = 84
5x = 75
x = 15
first number = x = 15
second number = 4x + 9 = 69